# Approximate Bayesian Computation sequential Monte Carlo via random forests

### v1

@@ 5.3 Stochastic biochemical reaction systems
Figure 5: Parameter inference for the Michaelis-Menten reaction system. a, b, c: Marginal distributions for reaction rates $c_{1}$ (a), $c_{2}$ (b) and $c_{3}$ (c), inferred with ABC-DRF and ABC-SMC-DRF, compared against true values (black lines). d: Absolute mean ($\mu^{*}$) and standard deviation ($\sigma$) of elementary effects of $c_{1},c_{2}$ and $c_{3}$, using Morris’s global sensitivity analysis. e, f, g, h: Range of molecule numbers for $E$ (e), $S$ (f), $ES$ (g) and $P$ (h) across time, simulated with $\theta\sim\pi_{ABC-DRF}\left(\theta\middle|s^{obs}\right)$ and $\pi_{ABC-SMC-DRF}\left(\theta\middle|s^{obs}\right)$ (box plots), compared against observed data $s^{obs}$ (black dots and lines).
@@ 6 Discussion
In this paper, we introduce a new Bayesian inference method, Approximate Bayesian Computation sequential Monte Carlo with random forests (ABC-SMC-(D)RF). It inherits random forest’s non-parametric nature, resulting in less dependence on user-defined hyperparameters compared to traditional ABC methods such as ABC-REJ, MCMC and ABC-SMC. The random forest is embedded in a sequential Monte Carlo regime which progressively updates the parameter distributions to focus on regions in the parameter space with higher likelihood. We perform numerical experiments for deterministic and stochastic models in ecology, population genetics and systems biology, and observe that ABC-SMC-(D)RF results in better posterior approximations than previous RF methods, and comparable results with traditional ABC implementations with optimized hyperparameters.
However, there are several areas of improvement that can extend ABC-SMC-(D)RF’s capabilities. First, a hyperparameter in ABC-SMC-(D)RF that potentially has an impact on the results is the choice of perturbation kernels $K_{t}\left(\theta\middle|\theta^{\prime}\right)$, which should balance between exploring the parameter space and targeting the regions already found to contain high likelihood from previous iterations. In ABC-MCMC and ABC-SMC, it is common to choose uniform or Gaussian kernels, but the optimal kernel form and parameterization for specific problems may be complicated (Drovandi and Pettitt, 2011; Lee, 2012). A typical approach is to employ kernels that are adjusted dynamically depending on the performance of the previous iteration (Beaumont et al., 2009; Del Moral et al., 2012; Liu et al., 2000; Filippi et al., 2013; Atchadé and Fort, 2010). A potentially promising improvement for ABC-SMC-(D)RF would be to adopt a similar adaptive kernel framework.
Second, it is sometimes not necessary to continue until the final iteration in ABC-SMC-(D)RF, if further computation is unlikely to yield significant improvements in the posterior distributions. There have been a number of papers determining the stopping criterion in ABC-SMC (for example, Prangle (2017)), and their incorporation in ABC-SMC-(D)RF promises to lower the computational cost while retaining accuracy.
Finally, in this paper we have focused on parameter inference. ABC-RF can perform model selection Raynal et al. (2019), and the same framework could potentially be implemented in ABC-DRF and ABC-SMC-(D)RF. ABC-SMC-(D)RF is designed as a wrapper around abcrf (Marin et al., 2022) and drf (Michel and Ćevid, 2021) and can therefore be updated whenever the original libraries are updated.
@@
The code for ABC-SMC-(D)RF and studies performed in this paper are available at https://github.com/dinhngockhanh/abc-smc-rf.
@@ Acknowledgments
The authors acknowledge the support from the Herbert and Florence Irving Institute for Cancer Dynamics and Department of Statistics at Columbia University.
@@
Appendix A Review of some ABC approaches
Assume we have a model $y=\mathcal{M}(\theta)$ that describes data $y$ with parameters $\theta\in\Theta$, and a prior distribution $\pi(\theta)$ that describes our a priori knowledge about the parameters. The posterior distribution of $\theta$ given observed data $y^{obs}$ is
$\pi(\theta|y^{obs})\propto f(y^{obs}|\theta)\cdot\pi(\theta),$
where $f(y^{obs}|\theta)$ is the likelihood function.
Approximate Bayesian Computation (ABC) was proposed as a practical method to approximate posterior distributions for problems in which the likelihood function is either theoretically intractable or computationally expensive. The data $y^{obs}$ is summarized with some low-dimensional statistic $s^{obs}=S(y^{obs})$. ABC methods then find the posterior distribution $\pi\left(\theta\mid d(s,s^{obs})<\epsilon\right)$, where $\epsilon$ is a tolerance threshold and $d(s,s^{obs})$ is a metric to compare $s^{obs}$ and statistics $s$ from data simulated with parameter $\theta$. If the statistic $S$ is sufficient, then $\pi\left(\theta\middle|d(s,s^{obs})<\epsilon\right)$ approximates the likelihood-based $\pi\left(\theta\middle|y^{obs}\right)$ as $\epsilon$ approaches 0. Here we briefly review some existing methods that will be used in comparisons with the proposed random forest approaches.
A.1 Rejection method
The simplest implementation of the ABC approach is the rejection method (ABC-REJ) (Algorithm 7), originally developed in population genetics [Tavaré et al., 1997, Fu and Li, 1997, Pritchard et al., 1999, Beaumont et al., 2002]. The statistics $s^{*}$ from data $y^{*}$ simulated with parameters $\theta^{*}\sim\pi(\theta)$ are compared to the observed statistics $s^{obs}$ with a distance function $d$, and only $\theta^{*}$ with $d(s^{*},s^{obs})$ smaller than a tolerance threshold $\epsilon$ are retained. The accepted values $\left\{\theta^{(1)},\dots,\theta^{(N)}\right\}$ are taken to approximate the required posterior distribution.
1 for $i=1,\dots,N$ do
2       Sample $\theta^{*}\sim\pi(\theta)$
3      Simulate data $y^{*}$ from the model with parameter $\theta^{*}$
4      Compute summary statistics $s^{*}=S(y^{*})$
5
6Order $\theta^{*}$ by increasing $d\left(s^{*},s^{obs}\right)$
Select $\theta^{(1)},...$ from the top percentiles of $\left\{\theta^{*}\right\}$
Algorithm 7 ABC-REJ
A.2 Markov Chain Monte Carlo
For problems where the prior distribution and posterior distributions are significantly different, the ABC-REJ method might require sampling $\theta^{*}$ and simulating $y^{*}$ many times before finding a parameter set that satisfies $d(s^{*},s^{obs})<\epsilon$ Different techniques have been proposed to improve the acceptance rate and decrease the number of model simulations.
Markov chain Monte Carlo (MCMC) approximates the posterior distribution by constructing a Markov chain, where the conditional distribution of the next proposed parameters only depends on the current set [Metropolis et al., 1953, Hastings, 1970]. Under suitable regularity conditions, the Markov chain converges to the posterior distribution[Tierney, 1994]. The Metropolis-Hastings algorithm 8 builds a Markov chain based on the likelihood function, starting from an initial $\theta^{(0)}$ sampled from the prior distribution [Metropolis et al., 1953, Hastings, 1970]. At each step, it proposes the next parameter set from a Markov kernel function $q\left(\theta\middle|\theta^{\prime}\right)$. The resulting chain $\left\{\theta^{(1)},\dots,\theta^{(N)}\right\}$ approximates the true posterior distribution $\pi\left(\theta\middle|y^{obs}\right)$.
1 Sample $\theta^{(0)}\sim\pi(\theta)$
2for $i=1,\dots,N$ do
3       Propose candidate parameter $\theta^{*}\sim q\left(\theta\middle|\theta^{(i-1)}\right)$
4      Calculate the acceptance probability $p_{accept}=\min\left(1,\frac{f\left(\theta^{*}\middle|y^{obs}\right)\cdot\pi\left(\theta^{*}\right)\cdot q\left(\theta^{(i-1)}\middle|\theta^{*}\right)}{f\left(\theta^{(i-1)}\middle|y^{obs}\right)\cdot\pi\left(\theta^{(i-1)}\right)\cdot q\left(\theta^{*}\middle|\theta^{(i-1)}\right)}\right)$ and random threshold $p\sim$ Uniform$(0,1)$
5      If $p_{accept}>p$, then $\theta^{(i)}\leftarrow\theta^{*}$, otherwise $\theta^{(i)}\leftarrow\theta^{(i-1)}$
Algorithm 8 Metropolis-Hastings MCMC
Marjoram et al. [2003] developed an ABC-MCMC method (Algorithm 9), a likelihood-free alternative to the Metropolis-Hastings algorithm. In lieu of $f\left(\theta\middle|y^{obs}\right)$, a proposed $\theta^{*}$ is only accepted if $d(s^{*},s^{obs})<\epsilon$. Because $\theta^{*}$ is drawn close to a previously accepted parameter set, the condition is typically more likely to be satisfied compared to the ABC-REJ method [Sisson et al., 2018]. The Markov chain results in the ABC-MCMC posterior distribution $\pi_{ABC-MCMC}\left(\theta\mid d(s^{*},s^{obs})<\epsilon\right)$.
1 Sample $\theta^{(0)}\sim\pi(\theta)$
2for $i=1,\dots,N$ do
3       Propose candidate parameter $\theta^{*}\sim q\left(\theta\middle|\theta^{(i-1)}\right)$
4      Simulate data $y^{*}$ with $\theta^{*}$, and compute summary statistics $s^{*}=S(y^{*})$
5      If $d(s^{*},s^{obs})\geq\epsilon$, then $\theta^{(i)}\leftarrow\theta^{(i-1)}$ and return to Step 3 with $i\leftarrow i+1$
6      Calculate probability to accept $p_{accept}=\min\left(1,\frac{\pi\left(\theta^{*}\right)\cdot q\left(\theta^{(i-1)}\middle|\theta^{*}\right)}{\pi\left(\theta^{(i-1)}\right)\cdot q\left(\theta^{*}\middle|\theta^{(i-1)}\right)}\right)$ and random threshold $p\sim$ Uniform$(0,1)$
7      If $p_{accept}>p$, then $\theta^{(i)}\leftarrow\theta^{*}$, otherwise $\theta^{(i)}\leftarrow\theta^{(i-1)}$
Algorithm 9 ABC-MCMC
To improve reliability and accuracy of the results, it is common practice to discard the particles in the initial “burn-in” phase of the Markov chain in both MCMC and ABC-MCMC, as these particles are dependent on the starting point and need not be representative of the posterior distribution.
A.3 Sequential Monte Carlo
Because the accepted particles in ABC-MCMC are correlated, the Markov chain can be stuck in regions of the parameter space with low probability. If the tolerance threshold $\epsilon$ is too small, the chain might take a long time to collect enough accepted particles [Sisson et al., 2007]. Moreover, the Markov chain cannot easily be computed in parallel.
Toni et al. [2008] introduced the ABC sequential Monte Carlo (ABC-SMC) method (Algorithm 10) as a different approach to solve the problem of low acceptance rate in ABC-REJ. ABC-SMC refines the particle population drawn from the prior distribution through a series of ABC-REJ iterations with decreasing tolerances $\epsilon_{1}>\dots>\epsilon_{T}$, where the posterior distribution from iteration $t-1$ becomes the prior distribution for iteration $t$. To encourage particle dispersion, the particles for iterations $t>1$ are perturbed with a Markov kernel $K_{t}(\theta|\theta^{\prime})$. The final set of particles $\left\{\theta_{T}^{(1)},\dots,\theta_{T}^{(N)}\right\}$ from iteration $T$ approximates the ABC-SMC posterior distribution $\pi_{ABC-SMC}\left(\theta\middle|d\left(s^{*},s^{obs}\right)<\epsilon_{T}\right)$. Note that $\pi_{ABC-SMC}\left(\theta\middle|d\left(s^{*},s^{obs}\right)<\epsilon_{1}\right)$ corresponds to $\pi_{ABC-rej}\left(\theta\middle|d\left(s^{*},s^{obs}\right)<\epsilon_{1}\right)$ from the ABC-REJ method.
1 for $t=1,\ldots,T$ do
2       for $i=1,\dots,N$ do
3             if $t=1$ then
4                  Sample $\theta^{**}\sim\pi(\theta)$
5            else
6                   Sample $\theta^{*}$ from $\left\{\theta^{(i)}_{t-1}\right\}$ with weights $\left\{w^{(i)}_{t-1}\right\}$
7                  $\theta^{**}\sim K_{t}(\theta|\theta^{*})$
8                  If $\pi(\theta^{**})=0$, then return to Step 6
9
10            Simulate data $y^{**}$ with parameter $\theta^{**}$, and compute $s^{**}=S(y^{**})$
11            If $d(s^{**},s^{obs})\geq\epsilon_{t}$, then return to Step 3
12            Record $\theta_{t}^{(i)}\leftarrow\theta^{**}$ with weight $w_{t}^{(i)}=\begin{cases}1&\text{ if }t=0\\ \frac{\pi\left(\theta^{(i)}_{t}\right)}{\sum^{N}_{j-1}w_{t-1}^{(j)}\cdot K_{t}\left(\theta_{t-1}^{(j)}\middle|\theta_{t}^{(i)}\right)}&\text{ if }t>0\end{cases}$
13      Normalize the weights such that $\sum_{i=1}^{N}w_{t}^{(i)}=1$
Algorithm 10 ABC-SMC

### v2

@@ 5.2 Stochastic biochemical reaction systems
Figure 3: Parameter inference for the Michaelis-Menten reaction system. a, b, c: Marginal distributions for reaction rates $c_{1}$ (a), $c_{2}$ (b) and $c_{3}$ (c), inferred with ABC-DRF and ABC-SMC-DRF, compared against true values (black lines). d: Absolute mean ($\mu^{*}$) and standard deviation ($\sigma$) of elementary effects of $c_{1},c_{2}$ and $c_{3}$, using Morris’s global sensitivity analysis. e, f, g, h: Range of molecule numbers for $E$ (e), $S$ (f), ESES (g) and $P$ (h) across time, simulated with $\theta\sim\pi_{ABC-DRF}\left(\theta\middle|s^{\rm obs}\right)$ and $\pi_{ABC-SMC-DRF}\left(\theta\middle|s^{\rm obs}\right)$ (box plots), compared against observed data $s^{\rm obs}$ (black dots and lines).
5.3 Stochastic birth-death models
5.3.1 Linear birth-death process
We next examine the performance of ABC-SMC-DRF for the homogeneous linear birth-death process illustrated in Tavaré (2018), for which exact results may be found. The model starts at time 0 with $Z(0)=z_{0}$ individuals, each of which divides at rate $\lambda>0$ and dies at rate $\mu>0$, independently for each individual. The number of individuals at time $t>0$ is denoted $Z(t)$, with expected value $\mathbb{E}[Z(t)]=Z(0)\mathrm{e}^{(\lambda-\mu)t}$.
The probability $\mathbb{P}(Z(t)=j\mid Z(0)=i)$ that there are $j$ individuals at time $t$, given that there are $i$ individuals at time 0, follows from Keiding (1975) as
$\mathbb{P}(Z(t)=j\mid Z(0)=i)=\sum_{l=0}^{\min(i,j)}\binom{i}{l}(1-\alpha(t))^{l}\alpha(t)^{i-l}\cdot\binom{j-1}{i-l}(1-\beta(t))^{l}\beta(t)^{j-l},j=1,2,\ldots$ (10)
and
$\mathbb{P}(Z(t)=0\mid Z(0)=i)=\alpha(t)^{i},$ (11)
where $\alpha(t)$ and $\beta(t)$ are defined by
$\alpha(t)=\frac{\mu\left(e^{(\lambda-\mu)t}-1\right)}{\lambda e^{(\lambda-\mu)t}-\mu},\quad\beta(t)=\frac{\lambda}{\mu}\alpha(t),$  $\quad\text{ if }\lambda\neq\mu;$
$\alpha(t)=\beta(t)=\frac{\lambda t}{1+\lambda t},$  $\quad\text{ if }\lambda=\mu,$
(Kendall, 1948). The Markov property implies that the likelihood for $\lambda$ and $\mu$, given an observation $y^{\rm obs}=(Z(t_{1})=z_{1},\dots,Z(t_{n})=z_{n})$ with $z_{n}>0$, is
$f\left(\lambda,\mu\middle|y^{\rm obs}\right)=\frac{\prod_{i=0}^{n-1}\mathbb{P}(Z(t_{i+1})=z_{i+1}\mid Z(t_{i})=z_{i})}{\mathbb{P}(Z(t_{n})>0\mid Z(0)=z_{0})},$ (12)
which may be calculated from (10) and (11). The observations used in the example below are generated on a grid of time points $t_{1}=1/25,t_{2}=2/25,\ldots,t_{25}=1,$ and $z_{0}=10.$
Instead of approximating the posterior distribution of $\theta=(\lambda,\mu)$, we approximate that of the extinction parameter $\mu/\lambda$ and the Malthusian parameter $\lambda-\mu$, so that $\theta=(\mu/\lambda,\lambda-\mu)$. The following prior distributions were used:
$\lambda-\mu$ $\sim{\rm Gamma}({\rm shape}=3,{\rm rate}=1)$
$\mu/\lambda$ $\sim{\rm Beta}(5,5)$
These priors guarantee that $\lambda>\mu$, implying that the probability of extinction is less than 1. We evaluate ABC-SMC-RF and ABC-SMC-DRF with $T=4$ iterations, each with $N_{t}=5,000$ simulations in the reference table and the random forest in each iteration is constructed from $B=2,500$ trees. We implement the normal perturbation kernels $K_{t}(\lambda-\mu|\lambda^{*}-\mu^{*})=\mathcal{N}\left(\lambda^{*}-\mu^{*},2\cdot\sigma_{t-1}^{\lambda-\mu}\right)$ and $K_{t}\left(\mu/\lambda\middle|\mu^{*}/\lambda^{*}\right)=\mathcal{N}\left(\mu^{*}/\lambda^{*},2\cdot\sigma_{t-1}^{\mu/\lambda}\right)$, where $\sigma_{t-1}^{\lambda-\mu}$ and $\sigma_{t-1}^{\mu/\lambda}$ are the empirical variances of the Malthusian and extinction parameters from iteration t1t-1. This perturbation framework was proven to be efficient in the context of ABC-SMC (Beaumont et al., 2009). The values of $\sigma_{t-1}^{\lambda-\mu}$ and $\sigma_{t-1}^{\mu/\lambda}$ are reported in Table 2. Figure 4 present the joint and marginal posterior distributions from ABC-SMC-RF and ABC-SMC-DRF. The results from both methods are close to the true posterior distribution.
Variance ABC-SMC-RF iterations ABC-SMC-DRF iterations
$t=2$ $t=3$ $t=4$ $t=2$ $t=3$ $t=4$
$2\cdot\sigma_{t-1}^{\lambda-\mu}$ 0.1207 0.1166 0.1514 0.2349 0.2785 0.2187
$2\cdot\sigma_{t-1}^{\mu/\lambda}$ 0.0273 0.0323 0.0343 0.0231 0.0234 0.0209
Table 2: Empirical variances in ABC-SMC-RF and ABC-SMC-DRF across different iterations
Figure 4: Parameter inference for the linear bith-death branching process. a: Joint posterior distributions for the Malthusian parameter $\lambda-\mu$ and the extinction parameter $\mu/\lambda$ from ABC-SMC-DRF (yellow contours) against ground truth (Eq. 12, density heatmap in gray-scale). b, c: Marginal distributions for $\lambda-\mu$ (b) and $\mu/\lambda$ (c), inferred with ABC-SMC-RF and ABC-SMC-DRF.
5.3.2 Controlled branching process
We also investigate the application of ABC-SMC-(D)RF for a discrete-time controlled branching process developed by González et al. (2022). Let $Z_{n}$ be the population size at generation $n=0,1,\dots$, which is defined recursively via
$Z_{0}=1$
$Z_{n+1}=\sum_{j=1}^{\phi_{n}(Z_{n})}X_{n,j}$
where $\phi_{n}(z)\sim\mathrm{Bin}(\xi(z),\gamma)$ with $\xi(z)=z+\lfloor\log z\rfloor$ represents the number of the $z$ individuals in generation $n$ that will be parents of generation $n+1$, and $X_{n,j}$ is the number of offspring from the $j$th individual at generation $n$. González et al. (2022) modeled $X_{n,j}$ as i.i.d. with respect to $n$ and $j$, with probabilities $\boldsymbol{p}(\kappa)=\{p_{k}=\mathbb{P}(X_{n,j}=k),k=0,1,\dots,\kappa\}$ where $\kappa$ is the maximum offspring capacity per individual. The data is simulated with $\gamma=0.8$, $\kappa=4$ and $\boldsymbol{p}(\kappa)\sim\mathrm{Bin}(4,0.9)$. From data statistics $\left\{Z_{1},\dots,Z_{10}\right\}$ and $\phi_{9}(Z_{9})$ (e.g., the number of parents of the last generation), we seek to find parameters $\theta=\left\{\gamma,\kappa,\boldsymbol{p}(\kappa)\right\}$ from prior distributions
$\gamma$ $\sim\beta(1,1)$
$\kappa$ $\sim\mathcal{U}(2,15)$
$\boldsymbol{p}(\kappa)$ $\sim\mathcal{D}(\kappa+1,\alpha_{\kappa}=(1,\dots,1)).$
As in (González et al., 2022), we analyze the results by examining the mean offspring count $m=\sum_{k=1}^{\kappa}k\cdot p_{k}(\kappa)$ instead of $\boldsymbol{p}(\kappa)$.
Similar to the linear birth-death process above, we perform ABC-SMC-RF and ABC-SMC-DRF with $T=4$ iterations, $N_{t}=5,000$ simulations per iteration and $B=2,500$ trees in each random forest. We implement the uniform perturbations $K_{t}(\kappa|\kappa^{*})=\mathcal{U}(\kappa^{*}-3,\kappa^{*}+3)$, $K_{t}(\gamma|\gamma^{*})=\mathcal{U}(\gamma^{*}-0.05,\gamma^{*}+0.05)$ and $K_{t}(p_{j}|p_{j}^{*})=\mathcal{U}(p_{j}^{*}-0.05,p_{j}^{*}+0.05)$. Figure 5 shows the comparisons between ABC-SMC-(D)RF posterior distributions and the true parameter values $\kappa=4$, $m=3.6$ and $\gamma=0.8$.
Figure 5 shows the inferred marginal posterior distributions of $\kappa$ (Figure 5b), $m$ (Figure 5c), $\gamma$ (Figure 5d), as well as the joint posterior distribution of $(\gamma,m)$ from ABC-SMC-DRF (Figure 5a). The marginal distributions for $m$ and $\gamma$ from both ABC-SMC-RF and ABC-SMC-DRF are centered around the true values. However, the ABC-SMC-RF distribution for $m$ is more concentrated compared to ABC-SMC-DRF, and the reverse holds for $\gamma$. Table 3 also shows that the mean and median of each parameter distribution converge to the true values after successive ABC-SMC-DRF iterations, with decreasing variances. This results in good approximations for $m$ and $\gamma$, with relative errors $10\%$ for both parameters in the final iteration.
Both ABC-SMC-RF and ABC-SMC-DRF overestimate $\kappa$ (Figure 5b). However, these distributions are similar to the results reported by González et al. (2022) (Figure 1 in (González et al., 2022)). As noted by the authors, the model with maximum offspring count $\kappa$ behaves identically to one with $\kappa^{\prime}>\kappa$ where $p_{\kappa+1}=\dots=p_{\kappa^{\prime}}=0$. Therefore, $\kappa$ can be inferred to be greater than the true value $\bar{\kappa}$, as long as the probabilities $p_{j}$ with $j>\bar{\kappa}$ are negligible (e.g., $10^{-3}$). We note that (González et al., 2022) utilizes a multi-step inference framework, where $\kappa$ is estimated first with ABC-SMC and $(\gamma,m)$ are then inferred with ABC-rejection, conditioned on the selected $\kappa$. In comparison, our framework of using ABC-SMC-(D)RF to infer $\kappa$, $m$ and $\gamma$ simultaneously may be more practical for implementation.
There are multiple sources of variability in the ABC-SMC-(D)RF results, including both the stochastic simulations that form the reference table in each iteration and the construction of the random forests. Therefore, we perform the ABC-SMC-(D)RF inference for the controlled branching process 500 times and analyze the changes in the inferred distributions (Table 4). Overall, there is some evidence that ABC-SMC-DRF underestimates $m$ and overestimates $\gamma$; the true values of both $m=3.6$ and $\gamma=0.8$ fall outside of their respective 95% interquantile ranges, $q_{2.5\%},q_{97.5\%}]$. In comparison, the interquantile ranges from ABC-SMC-RF successfully contain the true values, albeit with higher variance for $m$. This agrees with the results from the single inference run shown in Figure 5.
Statistics ABC-SMC-DRF iterations
1 2 3 4 (Final posterior)
$\mathbb{E}(\kappa)$ 6.326 5.478 5.239 5.406
$q_{50\%}(\kappa)$ 6 5 5 5
Var($\kappa$) 3.297 1.900 1.332 1.076
$\mathbb{E}(\gamma)$ 0.833 0.856 0.854 0.823
$q_{50\%}(\gamma)$ 0.873 0.872 0.869 0.825
Var($\gamma$) 0.013 0.007 0.006 0.005
$\mathbb{E}(m)$ 3.368 3.225 3.258 3.353
$q_{50\%}(m)$ 3.106 3.137 3.165 3.332
Var($m$) 0.500 0.243 0.234 0.259
Table 3: Means, medians and variances of the marginal posterior distributions for the controlled branching process from ABC-SMC-DRF.
Method $\kappa$ $\gamma$ $m$
$\mathbb{E}$ $q_{2.5\%},q_{97.5\%}]$ std $\mathbb{E}$ $q_{2.5\%},q_{97.5\%}]$ std $\mathbb{E}$ $q_{2.5\%},q_{97.5\%}]$ std
ABC-SMC-RF 5.8 [5, 7] 0.6 0.817 [0.795, 0.841] 0.012 4.03 [3.05, 5.05] 0.53
ABC-SMC-DRF 5.3 [5, 6] 0.5 0.864 [0.832, 0.902] 0.020 3.34 [3.15, 3.56] 0.10
Table 4: Mean ($\mathbb{E}$), $2.5^{th}$ and $97.5^{th}$ percentiles ($q_{2.5\%},q_{97.5\%}]$), and standard deviation (std) of ABC-SMC-(D)RF posterior marginal distribution medians for the controlled branching process across 500 runs. The data statistics are the same across different inference runs.
Figure 5: Parameter inference for the controlled branching process. a: Joint posterior distributions for $m$ and $\gamma$ from ABC-SMC-DRF (yellow contours) against true values (broken lines). b, c, d: Marginal distributions for $\kappa$ (b), $m$ (c) and $\gamma$ (d), inferred with ABC-SMC-RF and ABC-SMC-DRF, compared to true values (black lines).
5.3.3 A logistic branching process
Our final branching example concerns the logistic branching process discussed by Lambert (Lambert, 2005). This has state space $\{0,1,2,\ldots\}$ and non-zero transition rates given by
$n\to n+1$ $\mathrm{at\ rate\ }n\lambda;$
$n\to n-1$ $\mathrm{at\ rate\ }n(\mu+(n-1)c),$ (13)
for non-negative parameters $\lambda,\mu,c.$ The case $c=0$ corresponds to the model discussed in Section 5.3.1, and we assume henceforth that $c>0$. Lambert showed that there are two dichotomous behaviors in this model, depending on whether $\mu=0$ or $\mu>0$. When $\mu=0$ the population size $Z(\cdot)$ remains positive and converges in distribution to a Poisson random variable $P$ with mean $\lambda/c$, conditional on $P>0$. When $\mu>0$ the population goes extinct with probability 1. If we observe the population size at time points $t_{1}<t_{2}<\ldots<t_{n}$ and it happens that $Z(t_{n})>0$, we might ask whether there is evidence that $\mu=0$ or $\mu>0$.
This question can be addressed in principle using abcrf and drf approaches, by assuming priors for $\mu$ and $c$ (and, w.l.o.g., $\lambda=1$), generating a reference set and proceeding as earlier. One object of interest is the posterior probability that $\mu=0$, which gives information about the likelihood the population will survive.
Rather than detail the results here, we note that this example seems to be a case where the SMC version of the random forest methods is not feasible. This is because the algorithms successively sample new particles from the previous iteration and perturb them with Markov kernels, assuming smooth transitions. The perturbed particles in iterations $t\geq 2$ therefore include $\mu>0$ with probability one, hence ABC-SMC-(D)RF is incapable of accurately evaluating the probability that $\mu=0$.
@@ 6 Discussion
In this paper, we introduce a new Bayesian inference method, Approximate Bayesian Computation sequential Monte Carlo with random forests (ABC-SMC-(D)RF). It inherits random forest’s non-parametric nature, resulting in less dependence on user-defined arguments compared to traditional ABC methods such as ABC-REJ, MCMC and ABC-SMC.
The random forest is embedded in a sequential Monte Carlo regime that progressively updates the parameter distributions to focus on regions in the parameter space with higher likelihood. We perform numerical experiments for deterministic and stochastic models in ecology, population genetics and systems biology, and observe that ABC-SMC-(D)RF typically results in better posterior approximations than previous RF methods. The Lotka-Volterra example illustrates that generic ABC-SMC-RF produces results comparable to traditional ABC implementations with optimized arguments.
However, there are several areas of improvement that can extend ABC-SMC-(D)RF’s capabilities. First, a hyperparameter in ABC-SMC-(D)RF that potentially has an impact on the results is the choice of perturbation kernels $K_{t}\left(\theta\middle|\theta^{\prime}\right)$, which should balance between exploring the parameter space and targeting the regions already found to contain high likelihood from previous iterations. In ABC-MCMC and ABC-SMC, it is common to choose uniform or Gaussian kernels, but the optimal kernel form and parameterization for specific problems may be complicated (Drovandi and Pettitt, 2011; Lee, 2012). A typical approach is to employ kernels that are adjusted dynamically depending on the performance of the previous iteration (Beaumont et al., 2009; Del Moral et al., 2012; Liu et al., 2000; Filippi et al., 2013; Atchadé and Fort, 2010). By default, ABC-SMC-(D)RF implements Gaussian kernels $K_{t}\left(\theta\middle|\theta^{\prime}\right)=\text{Normal}\left(\theta^{\prime},2\cdot\text{var}\left(\theta_{t-1}\right)\right)$, proposed by Beaumont et al. (2009) for ABC-SMC. Further work might evaluate other adaptive kernel frameworks.
Second, it is sometimes not necessary to continue iteration in ABC-SMC-(D)RF, if further computation is unlikely to yield significant improvements in the posterior distributions. There have been a number of papers determining the stopping criterion in ABC-SMC (for example, Prangle (2017)), and their incorporation in ABC-SMC-(D)RF promises to lower the computational cost while retaining accuracy.
Finally, in this paper we have focused on parameter inference. ABC-RF can perform model selection (Pudlo et al., 2016), and the same framework could potentially be implemented in ABC-DRF and ABC-SMC-(D)RF. This will be relevant for problems similar to the logistic branching process in Section 5.3.3, where there is interest in model selection. ABC-SMC-(D)RF is designed as a wrapper around abcrf (Marin et al., 2022) and drf (Michel and Ćevid, 2021) and can therefore be updated together with the original libraries.
@@
The code for ABC-SMC-(D)RF and studies performed in this paper are available at https://github.com/dinhngockhanh/abcsmcrf.
@@ Acknowledgments
The authors acknowledge the support from the Herbert and Florence Irving Institute for Cancer Dynamics at Columbia University.
@@
Supplementary Materials
Appendix A Dependence of ABC-SMC-(D)RF on the prior distributions
We examine the influence of the weight recalibration in ABC-SMC-(D)RF (Step 15, Algorithm 6 and Step 18, Algorithm 7) on the posterior distributions, by comparing ABC-DRF and ABC-SMC-DRF for the hierarchical Normal mean model in Section A. This example presents a case study for whether ABC-SMC-DRF converges to the correct posterior distribution when the prior distribution is highly non-uniform. The model, statistics and true joint posterior distribution are given in Section .
In Fig. 6, we compare the results from drf with 20,000 particles, and abcsmcrf with 20,000 particles per iteration and 20 iterations in total using Gaussian perturbations with adaptive variances formulated in [Beaumont et al., 2009]. As discussed in the main text, ABC-DRF can already approximate the true posterior distribution well. Successive ABC-SMC-DRF posteriors essentially remain constant and in agreement with the true distribution. This demonstrates that the weight calibrations in ABC-SMC-(D)RF, adopted from ABC-SMC, maintain the influence of the prior distributions and particle perturbations on the posterior distributions, and that ABC-SMC-(D)RF becomes stable as the iterations proceed.
Figure 6: Inference of $\theta=(\theta_{1},\theta_{2})$ in the hierarchical model, with $\alpha=4,\beta=5$. Joint posterior distributions for $\theta_{1}$ and $\theta_{2}$, inferred from ABC-DRF with $N=20,000$ simulations (yellow contours) and ABC-SMC-DRF with $N_{t}=20,000$ simulations for iterations $t=1,\dots,20$ (red contours) against ground truth (density heatmap in gray-scale), with marginal distributions for each parameter from ABC-DRF (yellow histogram) and ABC-SMC-DRF (red histogram) against ground truth (black histogram).